# On blow-ups of sets with finite fractional variation

### v1

@@ Abstract.
Given $\alpha\in(0,1)$ and a set $E\subset\mathbb{R}^{N}$ with locally finite fractional $\alpha$-variation, we show that for almost every $x\in\mathbb{R}^{N}$ with respect to the $\alpha$-variation measure of $\mathbf{1}_{E}$, if $E$ admits a non-trivial tangent set at $x$ with locally finite integer perimeter, then $E$ also admits a tangent half-space oriented by the fractional unit normal of $E$ at $x$.
@@ 1.1. Setting
For further details on the operators (1.1) and (1.2) and on the formula (1.3), we refer the reader to [silhavy20]. Building on the integration-by-parts formula (1.3), in our previous works [brue-et-al22, comi-stefani19, comi-et-al22, comi-stefani22, comi-stefani23a, comi-stefani23b, comi-stefani24a, comi-stefani24b], together with Giovanni E. Comi and several collaborators, we developed a new theory of distributional fractional Sobolev and $BV$ spaces.
@@ 1.2. Fractional variation
Given $p\in[1,\infty]$ and a non-empty open set $\Omega\subset\mathbb{R}^{N}$, we say that $u\in BV^{\alpha,p}_{\rm loc}(\Omega)$ if $u\in L^{p}(\mathbb{R}^{N})$ and
@@ 1.3. Blow-up Theorem
    if $M\notin\{\emptyset,\mathbb{R}^{N}\}$ and $P(M)<\infty$, then $M=(m,\infty)$ for some $m\in\mathbb{R}$.
The first part of Theorem˜1.1 follows from [comi-stefani19]*Th. 5.8 and Prop. 5.9, while the second part is taken from [comi-stefani24b]*Th. 1.7. Theorem˜1.1 can be regarded as the fractional counterpart of De Giorgi’s Blow-up Theorem [maggi12]*Th. 15.5, which states that, if $\mathbf{1}_{E}\in BV_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}E$, the reduced boundary of $E$, then $\operatorname{Tan}(E,x)=\{H^{+}_{\nu_{E}(x)}(x)\}$, where
@@ 1.4. Main result
$\operatorname{Tan}(E,x)\cap BV_{\rm loc}^{\star}(\mathbb{R}^{N})\neq\emptyset\implies H_{\nu_{E}(x)}^{+}(x)\in\operatorname{Tan}(E,x).$ (1.5)
Some comments are now in order. First, as shown by the aforementioned examples in [comi-stefani24b]*Props. 1.13 and 1.15, the implication (1.5) fails if $\operatorname{Tan}(E,x)=\left\{\emptyset\right\}$ or $\operatorname{Tan}(E,x)=\{\mathbb{R}^{N}\}$, so we must assume that a non-trivial blow-up exists. Second, Theorem˜1.2 is, at present, the closest result to De Giorgi’s Blow-up Theorem that we are able to achieve; namely, for $|D^{\alpha}\mathbf{1}_{E}|$-almost every $x\in\mathscr{F}^{\alpha}E$, if $E$ admits a non-trivial tangent set of locally finite integer perimeter at $x$, then some tangent of $E$ at $x$ is the half-space oriented by $\nu_{E}^{\alpha}(x)$. Third, we do not know whether this property holds for every $x\in\mathscr{F}^{\alpha}E$. In fact, it is not clear whether the proof of property (iv) in Theorem˜1.1 can be extended to the more general case $\mathbf{1}_{M}\in BV_{\rm loc}(\mathbb{R}^{N})$, thereby yielding Theorem˜1.2 for every $x\in\mathscr{F}^{\alpha}E$.
The statement of Theorem˜1.2—and its proof—was inspired by [ambrosio-et-al09]*Th. 1.2, where a similar result is obtained for sets of locally finite perimeter in Carnot groups. The idea underlying the proof of Theorem˜1.2, as well as that of the main result in [ambrosio-et-al09], is to exploit the principle—introduced by Preiss in [preiss87]*Th. 2.12 (see also [mattila95]*Th. 14.16)—that tangent measures to tangent measures are tangent measures (see Theorem˜2.1 below). Therefore, to obtain the conclusion of Theorem˜1.2, it suffices to ensure that at least one iterated tangent set of $E$ is a half-space, which is indeed the case whenever some non-trivial tangent set of $E$ has locally finite integer perimeter.
@@ 2.1. Radon measure
Given a non-empty set $X\subset\mathbb{R}^{N}$, we let $\mathscr{M}(X)$ and $\mathscr{M}_{\rm loc}(X)$ be the spaces of finite and locally finite signed Radon measures on $X$, respectively.
@@
$\lim_{k\to\infty}\int_{X}\varphi\,\mathrm{d}\mu_{k}=\int_{X}\varphi\,\mathrm{d}\mu\quad\text{for every }\varphi\in C_{c}(X).$ (2.1)
In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(X)$ as $k\to\infty$.
For vector-valued measures, we write $\mathscr{M}(X;\mathbb{R}^{m})$ and $\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$, with $m\in\mathbb{N}$. The notion of (local) weak convergence applies to $\mathbb{R}^{m}$-valued Radon measures by requiring (2.1) to hold componentwise. In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$ as $k\to\infty$. Accordingly, the total variation of $\mu\in\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$ on an open set $A\subset X$ is defined as
$|\mu|(A)=\sup\left\{\int_{X}\varphi\cdot\,\mathrm{d}\mu:\varphi\in C_{c}(X;\mathbb{R}^{m}),\ \operatorname{supp}\varphi\subset A,\ \|\varphi\|_{L^{\infty}}\leq 1\right\}.$
We recall that $|\mu|\in\mathscr{M}_{\rm loc}(X)$ for every $\mu\in\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$; see [maggi12]*Lem. 4.17. For a detailed presentation of the theory of Radon measures, we refer to [ambrosio-et-al00, falconer97, maggi12, mattila95].
@@ 3.1. Properties of tangent sets
showing that $D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)$. In addition, we must have $D^{\alpha}\mathbf{1}_{F}\neq 0$, since otherwise $\mathbf{1}_{F}$ would be constant by [comi-stefani24b]*Prop. 1.8, and thus $F\in\left\{\emptyset,\mathbb{R}^{N}\right\}$, a contradiction.
@@
Since $x\in\mathscr{F}^{\alpha}E$, possibly passing to a subsequence (which we do not relabel), by Theorem˜1.1 there exists $G\in\operatorname{Tan}(E,x)$ such that $\mathbf{1}_{G}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $E_{k}\to G$ in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ as $k\to\infty$. Again by [comi-stefani24b]*Th. 1.6(i), possibly passing to a further subsequence (which we do not relabel), we also have
@@ 3.2. Iterated tangent sets
Let $x\in\mathscr{F}^{\alpha}E$ be such that Corollary˜3.1(ii) holds. Then, for every $F\in\operatorname{Tan}(E,x)$ and every $y\in\operatorname{supp}\mathscr{F}^{\alpha}F$, we have $\operatorname{Tan}(F,y)\subset\operatorname{Tan}(E,x)$. By the definition in (3.1), we infer that $\operatorname{Tan}^{2}(E,x)\subset\operatorname{Tan}^{1}(E,x)$, and the conclusion follows by iteration. ∎
@@
Let $x\in\mathscr{F}^{\alpha}E$ be such that Theorem˜3.3 holds. By assumption, there exists $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$ such that $\mathbf{1}_{F}\in BV_{\rm loc}(\mathbb{R}^{N})$. Hence, we have $\operatorname{Tan}(F,y)=\{H_{\nu_{F}(y)}^{+}(y)\}$ for every $y\in\mathscr{F}F$. Since $\mathscr{F}F\subset\mathscr{F}^{\alpha}F$ by [comi-stefani24b]*Th. 1.12(iii), from the definition in (3.1) we deduce that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}^{2}(E,x)$ for every $y\in\mathscr{F}F$. This, in turn, by Theorem˜3.3, implies that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}(E,x)$ for every $y\in\mathscr{F}F$. Combining [comi-stefani24b]*Prop. 1.13 with Theorem˜1.1, we then obtain that, for every $y\in\mathscr{F}F$,
@@
As a consequence, $\nu_{F}(y)=\nu^{\alpha}_{E}(x)$ for every $y\in\mathscr{F}F$, which implies that $F=H^{+}_{\nu^{\alpha}_{E}(x)}(x_{0})$ for some $x_{0}\in\mathbb{R}^{N}$ by [maggi12]*Prop. 15.15. Thus $H^{+}_{\nu^{\alpha}_{E}(x)}(x_{0})\in\operatorname{Tan}(E,x)$ for some $x_{0}\in\mathbb{R}^{N}$, and therefore $H^{+}_{\nu^{\alpha}_{E}(x)}(x)\in\operatorname{Tan}(E,x)$ as in (1.4), concluding the proof. ∎

### v2

@@ Abstract.
Given $\alpha\in(0,1)$ and a set $E\subset\mathbb{R}^{N}$ with locally finite fractional $\alpha$-variation, we show that for $|D^{\alpha}\mathbf{1}_{E}|$-a.e. $x$, every non-trivial tangent set of $E$ at $x$ with locally finite integer perimeter is a half-space oriented by the fractional inner unit normal of $E$ at $x$.
@@ 1.1. Setting
For further details on the operators (1.1) and (1.2) and on the formula (1.3), we refer the reader to [silhavy20]. Building on the integration-by-parts formula (1.3), in our previous works [brue-et-al22, comi-stefani19, comi-et-al22, comi-stefani22, comi-stefani23a, comi-stefani23b, comi-stefani24a, comi-stefani24b], together with Giovanni E. Comi and several collaborators, we developed a new theory of distributional fractional Sobolev and $BV$ spaces. The literature around the fractional gradient (1.1) is growing very rapidly in multiple directions, both theoretical and applicative. For a non-exhaustive list of contributions, we refer to [almi-et-al25, alicandro-et-al25, antil-et-al24, bellido-et-al20, bellido-et-al21, barrios-medina21, braides-et-al24, caponi-et-al25, carrero-et-al25, kreisbeck-schonberger22, liu-et-al24, liu-xiao22, lo-rodrigues23, schikorra-et-al17, schonberger24, shieh-spector15, shieh-spector18, spector19, spector20] and to the references therein.
@@ 1.2. Fractional variation
We continue the study of fractional $BV$ functions. Given $p\in[1,\infty]$ and an open set $\Omega\subset\mathbb{R}^{N}$, we say that $u\in BV^{\alpha,p}_{\rm loc}(\Omega)$ if $u\in L^{p}(\mathbb{R}^{N})$ and
@@ 1.3. Blow-up Theorem
    if $M\notin\{\emptyset,\mathbb{R}\}$ and $P(M)<\infty$, then $M=(m,\infty)$ for some $m\in\mathbb{R}$.
The first part of Theorem˜1.1 follows from [comi-stefani19]*Th. 5.8 and Prop. 5.9, while the second part is taken from [comi-stefani24b]*Th. 1.7. Theorem˜1.1 can be regarded as the fractional counterpart of De Giorgi’s Blow-up Theorem [maggi12]*Th. 15.5, which states that if $\mathbf{1}_{E}\in BV_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}E$, the reduced boundary of $E$, then $\operatorname{Tan}(E,x)=\{H^{+}_{\nu_{E}(x)}(x)\}$, where
@@ 1.4. Main result
Given $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$, for $|D^{\alpha}\mathbf{1}_{E}|$-a.e. $x\in\mathscr{F}^{\alpha}E$ the following holds. If $F\in\operatorname{Tan}(E,x)\cap BV_{\rm loc}^{\star}(\mathbb{R}^{N})$, then $F=H_{\nu_{E}^{\alpha}(x)}^{+}(y)$ for some $y\in\mathbb{R}^{N}$. In particular,
$\operatorname{Tan}(E,x)\cap BV_{\rm loc}^{\star}(\mathbb{R}^{N})\neq\emptyset\implies H_{\nu_{E}^{\alpha}(x)}^{+}(x)\in\operatorname{Tan}(E,x).$ (1.5)
Some comments are in order. First, as shown by the examples in [comi-stefani24b]*Props. 1.13 and 1.15, the implication (1.5) does not hold when all tangent sets are trivial, so the assumption that a non-trivial blow-up exists is necessary. Second, Theorem˜1.2 yields a rigidity property: for $|D^{\alpha}\mathbf{1}_{E}|$-almost every $x\in\mathscr{F}^{\alpha}E$, every non-trivial tangent set with locally finite integer perimeter is a half-space oriented by $\nu_{E}^{\alpha}(x)$. This provides the closest available analogue of De Giorgi’s Blow-up Theorem in the fractional setting. Third, we do not know whether this rigidity holds at every point of $\mathscr{F}^{\alpha}E$. In particular, extending property (iv) in Theorem˜1.1 to the broader case $\mathbf{1}_{M}\in BV_{\rm loc}(\mathbb{R})$ remains open.
Theorem˜1.2 was inspired by [ambrosio-et-al09]*Th. 1.2, where an analogous statement is established in Carnot groups. The key ingredient is the principle, due to Preiss [mattila95, preiss87], that tangent measures to tangent measures are tangent measures (see Theorem˜2.1 below). This fact makes it possible to reduce the proof to the situation in which a tangent of $E$ has locally finite integer perimeter, a case for which the conclusion follows from the classical theory.
A broader motivation for Theorem˜1.2 comes from the study of sets with locally finite distributional fractional perimeter. While De Giorgi’s theorem describes the infinitesimal structure of sets of finite perimeter, the fractional case lacks an analogous rectifiability theory and even the appropriate infinitesimal models are not fully understood. In this direction, Theorems˜1.2 and 3.3 suggest that certain rigidity features of the classical theory may persist under suitable assumptions on the existence of non-trivial tangent sets.
@@ 2.1. Radon measure
Given an open set $\Omega\subset\mathbb{R}^{N}$, we let $\mathscr{M}(\Omega)$ and $\mathscr{M}_{\rm loc}(\Omega)$ be the spaces of finite and locally finite signed Radon measures on $\Omega$, respectively.
@@
$\lim_{k\to\infty}\int_{\Omega}\varphi\,\mathrm{d}\mu_{k}=\int_{\Omega}\varphi\,\mathrm{d}\mu\quad\text{for every }\varphi\in C_{c}(\Omega).$ (2.1)
In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(\Omega)$ as $k\to\infty$.
For vector-valued measures, we write $\mathscr{M}(\Omega;\mathbb{R}^{m})$ and $\mathscr{M}_{\rm loc}(\Omega;\mathbb{R}^{m})$, with $m\in\mathbb{N}$. The notion of (local) weak convergence applies to $\mathbb{R}^{m}$-valued Radon measures by requiring (2.1) to hold componentwise. In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(\Omega;\mathbb{R}^{m})$ as $k\to\infty$. Accordingly, the total variation of $\mu\in\mathscr{M}_{\rm loc}(\Omega;\mathbb{R}^{m})$ on an open set $A\subset\Omega$ is defined as
$|\mu|(A)=\sup\left\{\int_{\Omega}\varphi\cdot\,\mathrm{d}\mu:\varphi\in C_{c}(\Omega;\mathbb{R}^{m}),\ \operatorname{supp}\varphi\subset A,\ \|\varphi\|_{L^{\infty}}\leq 1\right\}.$
We recall that $|\mu|\in\mathscr{M}_{\rm loc}(\Omega)$ for every $\mu\in\mathscr{M}_{\rm loc}(\Omega;\mathbb{R}^{m})$; see [maggi12]*Lem. 4.17. For a detailed presentation of the theory of Radon measures, we refer to [ambrosio-et-al00, falconer97, maggi12, mattila95].
@@ 3.1. Properties of tangent sets
showing that $D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)$. In addition, we must have $D^{\alpha}\mathbf{1}_{F}\neq 0$, since otherwise $\mathbf{1}_{F}$ would be constant by [comi-stefani24b]*Prop. 1.8, and thus $F\in\{\emptyset,\mathbb{R}^{N}\}$, a contradiction.
@@
Since $x\in\mathscr{F}^{\alpha}E$, possibly passing to a subsequence (which we do not relabel), by Theorem˜1.1 there exists $G\in\operatorname{Tan}(E,x)$ such that $\mathbf{1}_{G}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $E_{k}\to G$ in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ as $k\to\infty$. Again by [comi-stefani24b]*Th. 1.6(i), possibly passing to a further subsequence (which we do not relabel), we also have that
@@ 3.2. Iterated tangent sets
Let $x\in\mathscr{F}^{\alpha}E$ be such that Corollary˜3.1(ii) holds and let $F\in\operatorname{Tan}(E,x)$. If $F\in\{\emptyset,\mathbb{R}^{N}\}$, then $\mathscr{F}^{\alpha}F=\emptyset$ and there is nothing to prove. If instead $F\notin\{\emptyset,\mathbb{R}^{N}\}$, then for every $y\in\mathscr{F}^{\alpha}F$ we have that $\operatorname{Tan}(F,y)\subset\operatorname{Tan}(E,x)$. By the definition in (3.1), we infer that $\operatorname{Tan}^{2}(E,x)\subset\operatorname{Tan}^{1}(E,x)$, and the conclusion follows by iteration. ∎
@@
Let $x\in\mathscr{F}^{\alpha}E$ be such that Theorem˜3.3 holds. By assumption, there exists $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$ such that $\mathbf{1}_{F}\in BV_{\rm loc}(\mathbb{R}^{N})$. Hence, we have that $\operatorname{Tan}(F,y)=\{H_{\nu_{F}(y)}^{+}(y)\}$ for every $y\in\mathscr{F}F$. Since $\mathscr{F}F\subset\mathscr{F}^{\alpha}F$ by [comi-stefani24b]*Th. 1.12(iii), from the definition in (3.1) we deduce that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}^{2}(E,x)$ for every $y\in\mathscr{F}F$. This, in turn, by Theorem˜3.3, implies that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}(E,x)$ for every $y\in\mathscr{F}F$. Combining [comi-stefani24b]*Prop. 1.13 with Theorem˜1.1, we then obtain that, for every $y\in\mathscr{F}F$,
@@
As a consequence, $\nu_{F}(y)=\nu^{\alpha}_{E}(x)$ for every $y\in\mathscr{F}F$, which implies that $F=H^{+}_{\nu^{\alpha}_{E}(x)}(x_{0})$ for some $x_{0}\in\mathbb{R}^{N}$ by [maggi12]*Prop. 15.15, and (1.5) follows by [leonardi00]*Props. 2.1 and 2.2. ∎